OpenStudy (anonymous):

The center of the circle whose equation is (x + 2)² + (y - 3)² = 25 is

OpenStudy (mathmale):

compare your (x + 2)² + (y - 3)² = 25 to the general equation of a circle whose center is at (h,k) and radius is r: (x-h)^2 + (y-k)^2 = r^2. Determining the value of k is easy here; determining the value of h requires a bit more thought and work.